# GameVibe: A Multimodal Affective Game Corpus

### v1

@@ Methods
The dataset is split into five directories as depicted in Fig. 3. The “Stimuli” directory contains the 1-minute game videos organised into subdirectories by session and named according to the game in each video. The “Questionnaire” directory contains participants’ responses to the post-experiment questionnaire in CSV format, and a Python script to process the responses and create visualisations. The “Annotations” folder contains the annotation data returned from PAGAN and the Python scripts required to process them; this folder has several subdirectories of importance. In the “Raw” subdirectory, the raw annotations outputted by PAGAN for each session can be found. The “Raw Combined” subdirectory contains the raw data assembled into a single CSV for each task, and the ground truth signals for the QA test (see Methods). The “Processed” subdirectory contains the processed annotation data for the QA tasks and the Engagement task for all sessions in NumPy file format. These processed files consist of a Python dictionary containing the following hierarchy: data is sorted into sessions, which in turn are sorted into participants, and further sorted by game where the participants’ annotations can be found. The “Latent Extraction” subdirectory contains Python scripts to extract the latents from the video and audio found in the “Stimuli” directory. We also include examples of extracted latents for a 3-second time window that can be used for further analysis. Finally, the “Analysis” directory contains the scripts used to conduct the analysis in this paper, and includes a “Processed” subdirectory with annotation data with outliers removed (see Methods).
Figure 4: GameVibe’s post-processing pipeline, visualizing the transformed labels and the output files of each stage.
@@
Annotator Quality Assurance: We first illustrate the reliability of our annotators by computing their average QA scores across our visual and auditory QA tasks (see Methods). Since both QA tasks involve an annotation task where the ground truth (screen brightness and audio frequency respectively) is known in advance, we use the Signed Differential Agreement (SDA) metric [36] to measure similarity. We chose SDA as it has been proven effective in our previous QA study [23] and, due to its bounded nature between -1 and 1, is an intuitive metric for selecting a filter threshold. Across all participants and both QA tasks, the average SDA score was $0.43\pm 0.15$, meaning that annotators correctly annotated $71.5\%$ of all time windows among QA signals. If we used a QA filtering threshold value similar to previous work[23] (SDA = 0, or $50\%$ accuracy), only 3 annotators (P6 and P7 from Session 2, P17 from Session 4) would be removed due to poor QA performance. This verifies that almost all our annotators understood the annotation process and could accurately produce labels before moving on to the engagement task. For the purposes of the technical validation of this paper, however, no participants are removed through this QA filtering in order to assess all raw data included in the dataset.
@@
Inter-Annotator Agreement: We use DTW (see Methods) to create distance matrices for each video in each session, in order to calculate the inter-annotator agreement between annotators’ normalized engagement traces on the same stimuli. We average those DTW values between pairs of participants on a per-session basis, since the same participant with the same identifier (e.g. P1) annotated every video in the same session. We report the average DTW distance across all stimuli (30) per session in Table 1, using different participant identifiers because different participants annotated different sessions. We observe that there are differences between annotators; only in Session 1 annotators were more in agreement. However, annotators are rarely consistently disagreeing with all other annotators; the most obvious instances of this are P6 (with DTW distances over 10 with all other annotators in Session 2) and P17 (with DTW distances over 11 with three other annotators in Session 4). We note that both annotators performed poorly in the QA tasks as well, and would have been removed via an SDA cut-off on the two QA tasks (see above). In general, there is a statistically significant ($p<0.05$) negative Pearson correlation between average DTW distance of each participant’s engagement traces with all others in the same session and SDA score averaged from the two QA tasks ($r=-0.56$). This is a promising finding for the quality assurance offered by the additional QA tasks, as the closer the annotators are to the known ground truth in those two tasks (an SDA score near 1), the less likely they are to disagree with other annotators in the more challenging engagement annotation task.
Outliers per participant: We detect outliers on the engagement traces (see Methods) based on the KDE of all DTW matrices of all videos in the same session, and report results with two filters for outlier removal. When using the $90\%$ filter (see Methods), 60 out of 600 annotations are removed (15 per session), whilst with the $80\%$ filter 120 annotations are removed (30 per session). Interestingly, while P7 would be removed due to poor QA task performance, none of their 30 traces are removed as outliers for the 90% filter and only 2 traces are removed as outliers for the 80% filter. On the other hand, P17 has the most outliers of all 20 participants: 11 of their 30 traces are removed as outliers for the 90% filter and an extraordinary 16 out of 30 removed for the 80% filter. Other participants with many removed outliers are P4 (with 9 outliers for 80% filter) and P3, P13, P20 (each with 8 outliers for 80% filter). While average DTW distance was correlated with the average SDA score from the two QA tasks, the same metric is not significantly correlated ($p>0.05$) with the number of outliers removed: $r=-0.33$ for removed outliers per participant with the 90% filter and $r=-0.25$ for outliers with the 80% filter.
@@
Usage Notes
We provide annotation data in CSV format and the respective video stimuli in .mp4 format, both formats can be processed in most software packages or programming language. We provide Python files (.py) for data processing and extraction of feature representations are provided. The GameVibe_README.md file details the organisation of the dataset, explaining the structure, naming convention, and specific contents of each file.
@@
The dataset can be managed, visualised and pre-processed using Python files (.py) files. These files are accessible for download at: https://osf.io/p4ngx/?view_only=fc9c68cf3f104ad7afba5ab73a1c66a8.

### v2

@@ Methods
The dataset and files described in this section are publicly available and can be found on our Open Science Framework repository [40]. We provide annotation data in CSV format and the respective video stimuli in .mp4 format: both formats can be processed in most software packages or programming language. We provide Python files for data processing and feature extraction. The GameVibe_README.md file details the organisation of the dataset, explaining the structure, naming convention, and specific contents of each file.
The dataset is split into five directories as depicted in Fig. 4. The “Stimuli” directory contains the 1-minute game videos organised into subdirectories by session and named according to the game in each video. The “Questionnaire” directory contains participants’ responses to the post-experiment questionnaire in CSV format, and a Python script to process the responses and create visualisations. The “Annotations” directory contains the annotation data returned from PAGAN and the Python scripts required to process them; this folder has several subdirectories of importance. In the “Raw” subdirectory, the raw annotations outputted by PAGAN for each session can be found. The “Raw Combined” subdirectory contains the raw data assembled into a single CSV for each task, and the ground truth signals for the QA test (see Methods). The “Processed” subdirectory contains the processed annotation data for the QA tasks and the Engagement task for all sessions in NumPy file format. These processed files consist of a Python dictionary containing the following hierarchy: data is sorted into sessions, which in turn are sorted into participants, and further sorted by game where the participants’ annotations can be found. The “Latent Extraction” subdirectory contains Python scripts to extract the latents from the audiovisual data found in the “Stimuli” directory. We also include examples of extracted latents for a 3-second time window that can be used for further analysis. Finally, the “Analysis” directory contains the scripts used to conduct the analysis in this paper, and includes a “Processed” subdirectory with annotation data with outliers removed (see Methods).
Figure 5: GameVibe’s post-processing pipeline, visualising the transformed annotations and the output files of each stage.
@@
Annotator Quality Assurance: We first illustrate the reliability of our annotators by computing their average QA scores across our visual and auditory QA tasks (see Methods). Since both QA tasks involve an annotation task where the ground truth (screen brightness and audio frequency respectively) is known in advance, we use the Signed Differential Agreement (SDA) metric [39] to measure similarity. We chose SDA as it has been proven effective in our previous QA study [23] and, due to its bounded nature between -1 and 1, is an intuitive metric for selecting a filter threshold. Across all participants and both QA tasks, the average SDA score was $0.43\pm 0.15$, meaning that annotators correctly annotated $71.5\%$ of all time windows among QA signals. If we used a QA filtering threshold value similar to previous work[23] (SDA = 0, or $50\%$ accuracy), only 3 annotators (P6 and P7 from Session 2, P17 from Session 4) would be removed due to poor QA performance. This verifies that almost all our annotators understood the annotation process and could accurately produce annotations before moving on to the engagement task. For the purposes of the technical validation of this paper, however, no participants are removed through this QA filtering in order to assess all raw data included in the dataset.
@@
Inter-Annotator Agreement: We use DTW (see Methods) to create distance matrices for each video in each session, in order to calculate the inter-annotator agreement between annotators’ normalised engagement traces on the same stimuli. We average those DTW values between pairs of participants on a per-session basis, since the same participant with the same identifier (e.g. P1) annotated every video in the same session. We report the average DTW distance across all stimuli (30) per session in Table 1, using different participant identifiers because different participants annotated different sessions. We observe that there are differences between annotators; only annotators in Session 1 were more in agreement. However, annotators are rarely consistently disagreeing with all other annotators; the most obvious instances of this are P6 (with DTW distances over 10 with all other annotators in Session 2) and P17 (with DTW distances over 11 with three other annotators in Session 4). We note that both annotators performed poorly in the QA tasks as well, and would have been removed via an SDA cut-off on the two QA tasks (see above). In general, there is a statistically significant ($p<0.05$) negative Pearson correlation between average DTW distance of each participant’s engagement traces with all others in the same session and SDA score averaged from the two QA tasks ($r=-0.56$). This is a promising finding for the quality assurance offered by the additional QA tasks, as the closer the annotators are to the known ground truth in those two tasks (an SDA score near 1), the less likely they are to disagree with other annotators in the more challenging engagement annotation task.
Outliers per participant: We detect outliers on the engagement traces (see Methods) based on the KDE of all DTW matrices of all videos in the same session, and report results with two filters for outlier removal. When using the $90\%$ filter (see Methods), 60 out of 600 annotations are removed (15 per session), whilst with the $80\%$ filter 120 annotations are removed (30 per session). Interestingly, while P7 would be removed due to poor QA task performance, none of their 30 traces are removed as outliers for the 90% filter and only 2 traces are removed as outliers for the 80% filter. On the other hand, P17 has the most outliers of all 20 participants: 11 of their 30 traces are removed as outliers for the 90% filter and an extraordinary 16 out of 30 removed for the 80% filter. Other participants with many removed outliers are P4 (with 9 outliers for 80% filter) and P3, P13, P20 (each with 8 outliers for 80% filter). The average SDA score from the two QA tasks is not significantly correlated ($p>0.05$) with the number of outliers removed: $r=-0.33$ for removed outliers per participant with the 90% filter and $r=-0.25$ for outliers with the 80% filter.
@@
The data have been deposited into an Open Science Framework repository [40] and is available for public use. The dataset can be managed, visualised and pre-processed using Python files files.
Acknowledgements
This research is supported by the Foundation.AI (Foundation Modelling of Games via Artifcial Intelligence) research project funded by the Ministry for Education, Sport, Youth, Research and Innovation in Malta, as well as by the Malta Council for Science and Technology through the SINO-MALTA Fund 2022, Project OPtiMaL. Makantasis was supported by Project ERICA (GA: REP-2023-36) fnanced by the Xjenza Malta (Malta Council for Science & Technology - MCST), for and on behalf of the Foundation for Science and Technology,through the FUSION: R&I Research Excellence Programme.
Author contributions statement
G.N.Y conceived the experiments, M.B, M.K and K.P conducted the experiments under the supervision of K.M, A.L and G.N.Y, and M.B and M.K performed the analysis of results. All authors discussed the results, and all authors contributed equally to the final manuscript.
Competing interests
The authors declare no competing interests.